# Embedding in Recommender Systems: A Survey

### v2

@@ 3. Graph based Embedding
Recent advancements in graph machine learning have enabled off-the-shelf models to effectively capture complex relationships within graphs, producing node and graph embeddings that can be utilized for various downstream tasks. In recommender systems, users, items, attributes, and other auxiliary information can be used to create a variety of graph structures (e.g., social networks (He and Chu, 2010), user-item bipartite graphs (He et al., 2016), and knowledge graphs (Wang et al., 2018b)). Recently, graph-based embeddings in recommender systems have received a lot of attention. As graph embeddings can leverage the topology information in graph data, they are shown to be effective in various tasks in recommender systems. This section classifies methods for learning graph embeddings in recommender systems based on the categories of graph data structures: 1) Embedding of Homogeneous Graphs; 2) Embedding of Bipartite Graphs; 3) Embedding of Heterogeneous Graphs; 4) Embedding of Hypergraphs.
Figure 5. Graph types in the recommendation system
@@ 3.1. Embedding of Homogeneous Graphs
Where $h_{v_{i}}$ denotes the embedding of node $v_{i}$, $\theta$ denotes the parameters of the neural network. In this way, Deepwalk captures the structural information of the graph and generates similar embeddings for neighboring nodes. Asymmetric Proximity Preserving (APP) graph embedding (Zhou et al., 2017) points out that in many downstream recommender system applications, the nodes in the graph do not have symmetry. For example, the probability of observing a user purchasing a computer $i_{c}$ before buying a mouse $i_{m}$ is much higher than that of seeing a user buying a mouse $i_{m}$ before purchasing a computer $i_{c}$. Deepwalk is based on skip-gram needs to predict the context of each node and cannot capture the asymmetry of the nodes, and APP graph embedding only allows gradient updates in the direction of the sampled path. InfoWalk (Zhou et al., 2021) pointed out that there are nodes without in-degree in the directed graph, that is, dangling nodes, which cannot be handled by APP graph embedding because it cannot access such nodes.
However, as the graph size increases the number of possible different node sequences generated by random walk-based methods becomes not handleable. Some works try to solve this problem, such as LINE (Tang et al., 2015) and GNNs (e.g., GCN (Kipf and Welling, 2016), GAT (Veličković et al., 2017b)), which are widely used in recommender systems. GNNs do not need to sample node sequences, they can be directly applied to the whole graph. To learn high-quality graph embedding that can be applied to downstream tasks of recommender systems, GNNs need to be extended to take the special properties of recommender system tasks. HyperSoRec (Wang et al., 2021b) proposes that the user-user graph can be approximated as a structure of an $n$-order tree graph. Chami et al. (Chami et al., 2019) show that tree graph embedding in hyperbolic space can achieve better performance than its counterpart in Euclidean space which may suffer from severe distortion. HyperSoRec develops a hyperbolic mapping layer to map graph embedding in euclidean space to hyperbolic space. M2GRL (Wang et al., 2020b) proposes to combine homogeneous graphs of multiple views to enhance sparse features and enrich node information. It constructs catalog view, instance view and shop view, then adopts random walk and SGNS (Levy and Goldberg, 2014) to learn an independent embedding of each view, and finally, map them to a shared embedding space and align them separately. DG-ENN (Guo et al., 2021b) also uses more than one view of homogeneous graphs to learn embeddings jointly. It first learns embeddings from the user-user and item-item homogeneous graphs. Then it leverages the user-item graph as a regularization to ensure the learned embeddings are smooth.
@@ 3.2. Embedding of Bipartite Graphs
The bipartite graph is a unique structure with two node sets, where edges connect nodes from different sets. This is commonly used in recommendation systems (e.g. user-item interaction graphs), represented as $\mathcal{G}=\left(\mathcal{V}_{A}\cup\mathcal{V}_{B},\mathcal{E}\right)$, with $\mathcal{V}_{A}$ and $\mathcal{V}_{B}$ as node sets, and $\mathcal{E}$ denoting the edges.
Graph Convolutional Matrix Completion (GC-MC) (Berg et al., 2017) predicts ratings using a graph encoder to aggregate information from neighboring nodes with uniform weight. It aggregates user embeddings from one-hot embeddings based on interacted items, and item embeddings from one-hot vectors of interacting users. A bilinear decoder then reconstructs user-item ratings. However, GC-MC has drawbacks: 1) It employs one-hot encoding, hindering new node processing and becoming inefficient as the graph size grows. 2) GC-MC overlooks node features. STAR-GCN (Zhang et al., 2019b) resolves these by end-to-end learning of low-dimensional node embeddings. It simulates embedding new nodes with a masked vector, training the model to reconstruct embeddings. NGCF (Wang et al., 2019b) overcomes the second drawback of GC-MC by considering the node features, and it learns embeddings as follows:
(8) $\mathbf{n}_{u}^{(l)}=\sum_{i\in\mathcal{N}_{u}}\frac{1}{\sqrt{\left|\mathcal{N}_{u}\right|\left|\mathcal{N}_{i}\right|}}\left(\mathbf{W}_{1}^{(l)}\mathbf{h}_{i}^{(l)}+\mathbf{W}_{2}^{(l)}\left(\mathbf{h}_{i}^{(l)}\odot\mathbf{h}_{u}^{(l)}\right)\right)$
(9) $\mathbf{h}_{u}^{(l+1)}=\operatorname{LeakyReLU}\left(\mathbf{h}_{u}^{(l)}+\mathbf{n}_{u}^{(l)}\right)$
Where $\mathcal{N}_{u}$ and $\mathcal{N}_{i}$ denote the number of neighbor nodes of user $u$ and item $i$. The $\mathbf{h}_{u}^{(l)}$ and $\mathbf{h}_{i}^{(l)}$ denote the node embeddings of user $u$ and item $i$ in layer $l$. $W$ stands for linear transformation, which can be a fully connected layer.
LightGCN (He et al., 2020) argues that as the inputs of the user and item stem from ID embeddings lacking semantic information, there is no need for nonlinear transformations. Therefore, LightGCN introduces a straightforward and efficient aggregation method:
@@
Where $\mathcal{N}_{u}$ and $\mathcal{N}_{i}$ denote the number of neighbor nodes of user $u$ and item $i$.The $\mathbf{h}_{u}^{(l)}$ and $\mathbf{h}_{i}^{(l)}$ denote the node embeddings of user $u$ and item $i$ in layer $l$.
Although LightGCN greatly simplifies GCN, it still requires a long training time because multilayer message passing dominates the training of the model. The relatively slow training limits the application of LightGCN in real recommendation scenarios. UltraGCN (Mao et al., 2021) further removes the multi-layer messaging process by directly optimizing the cosine similarity of nodes and neighbors to capture the higher-order collaborative signals between users and items, and it uses a negative sampling strategy to avoid oversmoothing.
Recognizing the inherent sparsity within user-item interaction graphs, Collaborative Similarity Embedding (CSE) (Chen et al., 2019) suggests enhancing embeddings by incorporating higher-order similarities among nodes of the same type. CSE introduces two modules, namely DSEmbed and NSEmbed, to capture user-item similarity and inter-item/inter-user similarity within a user-item bipartite graph. DSEmbed employs an optimization scheme based on rating data, aiming to model the proximity between users and items. This is achieved by maximizing the log-likelihood function of positive and negative samples. The optimization objective of DSEmbed can be formulated as:
(11) $\arg\max_{\Phi}\sum_{\left(v_{i},v_{j},v_{k}\right)}\log p\left(v_{j}>_{i}v_{k}\mid\mathbf{h}\right)$
where $p\left(v_{j}>_{i}v_{k}\mid\mathbf{h}\right)$ is calculated as follows:
(12) $p\left(v_{j}>_{i}v_{k}\mid\mathbf{h}\right)=Sigmoid\left(\mathbf{h}_{i}^{T}\cdot\mathbf{h}_{j}-\mathbf{h}_{i}^{T}\cdot\mathbf{h}_{k}\right)$
Conversely, NSEmbed employs random walks and CBOW (Mikolov et al., 2013a) to predict central nodes, enabling the modeling of higher-order neighbor relationships among users or items. Another approach, GE (Xie et al., 2016), proposes joint training to address the sparsity observed in user-POI bipartite graphs. This is achieved by incorporating multiple graph perspectives, where supplementary bipartite graphs (e.g., POI-POI, POI-Region) augment the model’s efficacy by providing enriched information.

### v3

@@ 4. Graph based Embedding
Graph-based embeddings have gained significant attention in recommender systems due to their ability to leverage topological information in graph data. This capability has been greatly enhanced by recent advancements in graph machine learning, which enable models to effectively capture complex relationships and produce high-quality node and graph embeddings for various downstream tasks. In recommender systems, these techniques can be applied to a variety of graph structures, such as social networks (He and Chu, 2010), user-item bipartite graphs (He et al., 2016), and knowledge graphs (Wang et al., 2018b), which are created using users, items, attributes, and other auxiliary information. This section begins with an overview of graph representation learning techniques for embeddings, including spectral and spatial Graph Neural Networks (GNNs), graph pooling techniques, and self-supervised learning approaches for graph embedding. These methods form the foundation for various graph-based embedding approaches in recommender systems. Following this, we explore embedding methods based on various graph data structures: 1) Homogeneous Graphs; 2) Bipartite Graphs; 3) Heterogeneous Graphs; 4) Hypergraphs. Finally, we review relevant surveys and discuss future directions in this field. The techniques and models for each graph structure are summarized in Table 4.
Table 4. Graph-Based Embedding
Graph-Based Embedding Graph Representation Learning Techniques (§4.1) Spectral GNNs: SpectralCF (Zheng et al., 2018), JSCN (Liu et al., 2019), SComGNN (Luo et al., 2024)
Spatial GNNs: Message-Passing Mechanism (Kipf and Welling, 2016; Gilmer et al., 2017), Graph Transformers (Dwivedi and Bresson, 2020; Rampášek et al., 2022)
Graph Pooling: DiffPool (Ying et al., 2018), SR-GNN (Wu et al., 2019b), RGNN (Liu et al., 2021e), HG-Pool (Wu et al., 2021c)
Graph SSL: SGL (Wu et al., 2021b), HHGR (Zhang et al., 2021a), G-BERT (Shang et al., 2019), PT-GNN (Hao et al., 2021), BUIR (Lee et al., 2021)
Embedding of Homogeneous Graphs 4.2) Random Walk Models: Deepwalk (Perozzi et al., 2014), APP (Zhou et al., 2017), InfoWalk (Zhou et al., 2021)
GNN-based Models: HyperSoRec (Wang et al., 2021b), M2GRL (Wang et al., 2020b), DG-ENN (Guo et al., 2021b)
Embedding of Bipartite Graphs (§4.3) GCN-based Models: GC-MC (Berg et al., 2017), STAR-GCN (Zhang et al., 2019), NGCF (Wang et al., 2019b), LightGCN (He et al., 2020), UltraGCN (Mao et al., 2021)
Similarity-based Models: Collaborative Similarity Embedding (CSE) (Chen et al., 2019), GE (Xie et al., 2016)
Embedding of Heterogeneous Graphs 4.4) Social Recommendation: DiffNet (Wu et al., 2019a), GraphRec (Fan et al., 2019), DiffNet++ (Wu et al., 2020), DANSER (Wu et al., 2019c)
Knowledge Graph-based: KGCN (Wang et al., 2019c), KGAT (Wang et al., 2019a), KGIN (Wang et al., 2021a)
Embedding of Hypergraphs (§4.5) Key Models: IHGNN (Cheng et al., 2022), HyperGroup (Guo et al., 2021c), HEMR (La Gatta et al., 2022)
Figure 3. Graph types in the recommendation system. (a) Homogeneous Graph: all nodes are of the same type. (b) Bipartite Graph: A unique structure with two node sets, where edges connect nodes from different sets. (c) Heterogeneous Graph: Encompasses diverse node categories that can be interconnected. (d) Hypergraph: In a hypergraph, each hyperedge can connect arbitrarily many nodes.
4.1. Graph Representation Learning Techniques
4.1.1. Spectral GNNs
Spectral methods in graph theory are based on the eigendecomposition of graph matrices, particularly the graph Laplacian. These methods provide a theoretical foundation for understanding graph structures and have been extended to develop Spectral Graph Neural Networks (Spectral GNNs) (Bruna et al., 2013; Bo et al., 2023). Given an undirected graph $G$ with $n$ nodes, let $A$ be the adjacency matrix and $D$ be the diagonal degree matrix. The unnormalized graph Laplacian is defined as L=DAL=D-A and the normalized Laplacian is given by: $L_{sym}=I-D^{-1/2}AD^{-1/2}$. Spectral graph convolution is defined in the Fourier domain. Let $U$ be the matrix of eigenvectors of $L$, and $\Lambda$ be the diagonal matrix of its eigenvalues. The graph Fourier transform of a signal $x$ is defined as $F(x)=U^{T}x$, and the inverse transform as $F^{-1}(\hat{x})=U\hat{x}$. The spectral convolution of the signal $x$ with a filter $g_{\theta}$ is given by:
(6) $g_{\theta}*x=Ug_{\theta}(\Lambda)U^{T}x$
where $g_{\theta}(\Lambda)$ is a diagonal matrix with entries that are a function of the eigenvalues. In the context of recommender systems, this operation allows us to filter graph signals (e.g., user preferences) in the spectral domain, effectively capturing smooth variations across the user-item interaction graph. Recent advancements in Spectral GNNs have primarily focused on enhancing computational efficiency and analyzing expressive power (Bo et al., 2023). These theoretical developments have paved the way for practical applications in various domains, including recommender systems. In this context, spectral methods have demonstrated their effectiveness across different recommendation scenarios. SpectralCF (Zheng et al., 2018) applied spectral convolution to user-item bipartite graphs, improving cold-start recommendations. Extending to cross-domain scenarios, JSCN (Liu et al., 2019) proposed transferable spectral representations and joint spectral convolution on multiple graphs, incorporating a domain adaptive module to handle domain incompatibility. Further exploring the spectral properties of graphs, SComGNN (Luo et al., 2024) analyzed complementary item relationships in item graphs, identifying low-frequency and mid-frequency components in the spectral domain corresponding to relevance and dissimilarity attributes, respectively. This observation informed the design of spectral filters for modeling these relationships. Collectively, these studies illustrate the versatility of spectral methods in capturing complex interactions within recommendation tasks, highlighting their potential for future developments in the field of recommender systems.
4.1.2. Spatial GNNs
Spatial methods in Graph Neural Networks (GNNs) have gained significant popularity due to their intuitive design, scalability, and flexibility in handling diverse graph structures. These methods typically employ a message-passing mechanism, iteratively updating node representations by aggregating information from neighboring nodes (Kipf and Welling, 2016; Gilmer et al., 2017). The general update rule for node $i$ in the $k$-th layer can be expressed as:
(7) $\mathbf{x}_{i}^{(k)}=\operatorname{Aggregation}^{(k)}\left(\mathbf{x}_{i}^{(k-1)},\bigoplus_{j\in\mathcal{N}(i)}\operatorname{Message}^{(k)}\left(\mathbf{x}_{i}^{(k-1)},\mathbf{x}_{j}^{(k-1)},\mathbf{e}_{j,i}\right)\right),$
where $\mathbf{x}_{i}^{(k)}$ represents the updated node representation, $\mathcal{N}(i)$ denotes the neighborhood of node $i$, $\operatorname{Message}^{(k)}$ computes messages based on node and edge features, and $\operatorname{Aggregation}^{(k)}$ combines the current representation with aggregated messages. The operator $\bigoplus$ represents the aggregation function. A complete GNN typically stacks multiple such layers, denoted as $\mathrm{GNN}(\cdot)$ for brevity. Spatial methods capture and propagate local structural information across graphs, making them well-suited for recommender systems where user-item interactions are crucial. These methods have been widely adopted in various recommendation scenarios, including collaborative filtering (Wang et al., 2019b), session-based recommendations (Wu et al., 2019b), and social recommendations (Fan et al., 2019). Their flexibility allows the incorporation of heterogeneous data, enhancing the modeling of complex recommendation scenarios. Recent advances in spatial GNNs span both theoretical and architectural fronts. Theoretical work has focused on understanding GNN expressive power through the Weisfeiler-Leman (WL) hierarchy, proposing advanced methods to overcome $k$-WL limitations (Xu et al., 2018a; Wang and Zhang, 2024; Zhang et al., 2023a; Zhao et al., 2025). Architecturally, Graph Transformers have emerged as a powerful framework that combines message passing with global attention mechanisms (Dwivedi and Bresson, 2020; Rampášek et al., 2022; Kim et al., 2022; Zhao et al., 2023). In recommender systems, transformer-based graph models have been developed to capture long-range dependencies in user-item graphs (Xia et al., 2022) and model complex multi-hop relationships (Li et al., 2023b), effectively addressing the over-smoothing problem of deep GNNs while maintaining scalability for large-scale recommendation scenarios.
4.1.3. Graph Pooling
In recommender systems, graph pooling is an important technique for obtaining global abstract representations of complex relational structures, enabling the capture of structural patterns in user preferences, item characteristics, and their diverse interactions. This technique reduces graph complexity while preserving essential structural information, leading to more efficient and expressive recommendation models. Graph pooling enhances the ability to model multi-scale relationships and aggregate information across different granularities, which is particularly valuable for handling the heterogeneous and multi-faceted nature of recommendation data. Graph pooling typically aggregates node information to form more compact graph embeddings, emphasizing important structural patterns and relationships (Wang et al., 2024a). For example, DiffPool (Ying et al., 2018) learns soft assignments of nodes to clusters:
(8) $\mathbf{S}^{(l)}=\text{softmax}(\text{GNN}_{l,\text{pool}}(\mathbf{A}^{(l)},\mathbf{H}^{(l)}))$
where $\mathbf{S}^{(l)}$ is the assignment matrix, $\mathbf{A}^{(l)}$ is the adjacency matrix, and $\mathbf{H}^{(l)}$ is the node feature matrix at layer $l$. These pooling operations enhance GNNs’ ability to capture multi-scale structural information. In the field of graph learning, the development of more powerful pooling techniques that incorporate insights from graph theory (Bianchi et al., 2020; Tsitsulin et al., 2023) and topology (Ying et al., 2024), while not tailor-made for recommender systems, still offers a potential path forward. Graph pooling techniques have proven particularly valuable in recommender systems by enabling multi-scale relationship modeling. For capturing short-term preferences, SR-GNN (Wu et al., 2019b) employs pooling to aggregate item embeddings within sessions. Extending beyond session-based scenarios, RGNN (Liu et al., 2021e) introduces a personalized graph pooling operator to learn hierarchical representations from review graphs, thereby constructing semantic representations for both users and items. Most comprehensively, HG-Pool (Wu et al., 2021c) represents each user as a personalized heterogeneous graph and proposes a novel heterogeneous graph pooling method, allowing the model to capture diverse preference patterns across different entity types and relations.
4.1.4. Graph Embedding via Self-supervised learning
Self-supervised learning (SSL) has emerged as a powerful paradigm for leveraging unlabeled data in machine learning models. Building upon the foundations of spectral and spatial graph neural networks, Graph SSL (GSSL) techniques have shown promise in learning invariant and robust representations without extensive labeled data (Liu et al., 2022; Xie et al., 2022c). Several prominent GSSL frameworks have been proposed in recent years, each with its unique approach. GRACE (Zhu et al., 2020) uses edge perturbation and the InfoNCE estimator, while MVGRL (Hassani and Khasahmadi, 2020) introduces graph diffusion as augmentation and employs the Jensen-Shannon estimator as objective. GCC (Qiu et al., 2020) focuses on instance-level pretraining with subgraph sampling, and BGRL (Thakoor et al., 2021) adapts the BYOL approach to graphs without relying on negative samples. G-BT (Bielak et al., 2022) applies the Barlow Twins objective to learn invariant graph representations. These frameworks utilize augmentation techniques such as node/edge dropout, graph diffusion, and subgraph sampling to generate diverse graph views for SSL. Some recent work has explored GSSL based on spectral principles, focusing on how edge perturbation can be guided by the spectrum (Lin et al., 2023; Chen et al., 2024; Yang et al., 2024a; Jian et al., 2025). These generalized GSSL techniques show great promise in various graph-based tasks, and they open up new avenues for applications in recommender systems to exploit the rich structural information in various graphs. In the following, we explore GSSLs adapted and extended for recommendation tasks, specifically categorized into contrastive, generative, and predictive approaches (Yu et al., 2023). Each category offers unique advantages in capturing different aspects of graph structures to enhance recommendation performance.
Contrastive Learning Methods
Contrastive learning (CL) in recommendation systems aims to generate self-supervised signals by minimizing the distance between views of the same user or item and maximizing that between different users or items in the embedding space. In graph-based recommendations, CL methods typically involve three key components: data augmentation, a graph encoder, and CL loss. Common graph augmentation techniques in recommendation contexts, as illustrated in Figure 4, include: 1) Node/Edge Dropout: Removing user/item nodes or interaction edges to identify influential components and enhance system robustness. 2) Graph Diffusion: Introducing user-item similarity through new edges to model potential preferences. 3) Subgraph Sampling: Fabricating an augmented graph by selecting nodes and edges to accentuate local connectivity in the user-item graph.
Figure 4. Augmentation methods of Graph. Drop (a) interaction edges or (b) user/item nodes to identify influential components. (c) Introducing user-item similarity through new edges. (d) Sample local nodes and edges in the subgraph to accentuate connectivity.
The contrastive learning task can be formulated as:
(9) $f_{\theta}^{*}=\arg\min_{f_{\theta},g_{\phi}}\mathcal{L}_{cl}\left(g_{\phi}\left(f_{\theta}(\mathcal{V}_{1}),f_{\theta}(\mathcal{V}_{2})\right)\right)$
where $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$ are two views of the graph, $f_{\theta}$ is a GNN encoder, $g_{\phi}$ is a projection head, and $\mathcal{L}_{cl}$ is a contrastive loss function. Building upon traditional augmentation techniques, GraphAug (Zhang et al., 2024c) proposes a more principled approach to graph augmentation through information bottleneck regularization. Unlike conventional methods that rely on random perturbations, GraphAug automatically distills informative self-supervision information while filtering out noise, alleviating the fundamental challenge of noisy self-supervised signals in real-world recommendation environments. Several notable works have applied contrastive learning to recommendation tasks. SGL (Wu et al., 2021b) learns user and item embeddings in bipartite graphs using various augmentation techniques, while HHGR (Zhang et al., 2021a) proposes a dual-scale node-dropping strategy specifically for group recommendations. In the realm of cross-domain recommendations, CCDR (Xie et al., 2022a) and PCRec (Wang et al., 2021c) leverage contrastive learning to transfer knowledge between domains. DCL (Liu et al., 2021c) challenges conventional assumptions about negative samples by conducting subgraph sampling through perturbation of a given node’s L-hop ego network. Heterogeneous graphs, which contain diverse node and edge types common in complex recommender systems, have also been explored in recent contrastive learning works (Wang et al., 2021d; Zhu et al., 2022; Cai et al., 2022; Chen et al., 2023). These studies investigate how to leverage heterogeneous node and edge types to better reflect the multi-faceted nature of user preferences.
Generative Methods
Inspired by the Masked Language Model (MLM) paradigm, notably successful in models like BERT (Devlin et al., 2018), generative methods in this context often leverage the powerful ability of sequence models to learn embeddings by reconstructing corrupted data (e.g., sequences with masked items). BERT4Rec (Sun et al., 2019) utilizes the BERT (Devlin et al., 2018) encoder for sequential recommendation. In particular, it uses BERT to model sequential user behaviors. The generative SSL task can be formulated as:
(10) $f_{\theta}^{*}=\arg\min_{f_{\theta},g_{\phi}}\mathcal{L}_{gen}\left(g_{\phi}\left(f_{\theta}(\tilde{\mathcal{D}})\right),\mathcal{D}\right)$
where $\tilde{\mathcal{D}}$ is a corrupted version of the original data $\mathcal{D}$, and $\mathcal{L}_{gen}$ is typically a reconstruction loss such as cross-entropy or mean squared error. Applying MLM-style generative self-supervised learning directly to graph-structured data can be less straightforward than on purely sequential data. Consequently, many approaches in recommender systems first extract or transform graph data into sequences (e.g., user interaction sequences, node paths) and then apply sequence-based generative SSL techniques. For example, G-BERT (Shang et al., 2019) deals with the training data such that diagnosis and medical ontology are represented by a tree structure. The model utilizes the graph attention network (Veličković et al., 2017a) to learn the ontology embedding from diagnosis and medical tree-like graphs in order to make medication recommendations. Then, it feeds the ontology embeddings into a BERT-like model and pre-trains it with MLM-based tasks, which include the self-prediction task and the dual-prediction task. The former task is to predict the masked codes from the same type (i.e., medication or diagnosis) of ontology embeddings, and the latter is to reconstruct the masked codes from different types of ontology embeddings. PMGT (Liu et al., 2021d) focuses on multimodal recommendations, handling nodes with diverse information in a multimodal graph (e.g., item descriptions and images). They propose MCNSampling, a novel method to select neighboring nodes using importance scores and use BERT-like patterns for node embedding learning. PT-GNN (Hao et al., 2021), instead of reconstructing masked nodes or items, addresses cold-start nodes with few neighbors using a meta-learning approach. It first pre-trains a transformer encoder on the embeddings of the first $K$ neighboring nodes to simulate cold-start scenarios. Then, a GCN model is pre-trained, incorporating the meta embedding in each convolution step. It finally maximizes cosine similarity between the predicted embedding and the ground-truth embedding generated by NCF (He et al., 2017) using all neighbor nodes.
In conclusion, these generative methods leverage the successful MLM paradigm and the powerful modeling capabilities of Transformer encoders, aiming to learn high-performance and high-capacity embeddings. While MLM is naturally suited to sequential data, researchers have often adapted these techniques for graph embedding learning by transforming graph data into sequential representations.
Predictive Methods
Predictive SSL methods in recommendations generate new samples or labels from the original data to guide the learning process. These can be further divided into sample-based and pseudo-label-based approaches.
For pseudo-label-based methods, the task can be expressed as:
[…]
where $\hat{\mathcal{Y}}$ represents generated pseudo-labels, and $\mathcal{L}_{pred}$ is a prediction loss function.
The core of Predictive Methods is the construction of pseudo-labels, and CHEST (Wang et al., 2023b) proposes the use of meta-path type as a prediction target in heterogeneous user-item graphs scenarios. BUIR (Lee et al., 2021), inspired by BYOL (Grill et al., 2020), takes the representational consistency of the outputs of two encoders as a prediction task.
In conclusion, SSL techniques for graph embeddings extend the capabilities of spectral and spatial GNNs to scenarios with limited labeled data in recommender systems. By leveraging the unique graph structures inherent in recommender systems, these methods enhance the quality and robustness of recommendations. They capture complex relationships and implicit feedback, which are abundant in recommendation scenarios but difficult to model with traditional supervised approaches.
@@ 4.2. Embedding of Homogeneous Graphs
Where $h_{v_{i}}$ denotes the embedding of node $v_{i}$, $\theta$ denotes the parameters of the neural network. In this way, Deepwalk captures the structural information of the graph and generates similar embeddings for neighboring nodes. Asymmetric Proximity Preserving (APP) graph embedding (Zhou et al., 2017) points out that in many downstream recommender system applications, the nodes in the graph do not have symmetry. For example, the probability of observing a user purchasing a computer $i_{c}$ before buying a mouse $i_{m}$ is much higher than that of seeing a user buying a mouse $i_{m}$ before purchasing a computer $i_{c}$. Deepwalk is based on skip-gram and needs to predict the context of each node, and cannot capture the asymmetry of the nodes. APP graph embedding only allows gradient updates in the direction of the sampled path. InfoWalk (Zhou et al., 2021) pointed out that there are nodes without in-degree in the directed graph, that is, dangling nodes, which cannot be handled by APP graph embedding because it cannot access such nodes. However, as the graph size increases, the number of possible different node sequences generated by random walk-based methods becomes not handleable. Some works try to solve this problem, such as LINE (Tang et al., 2015) and GNNs (e.g., GCN (Kipf and Welling, 2016), GAT (Veličković et al., 2017b)), which are widely used in recommender systems. GNNs do not need to sample node sequences; they can be directly applied to the whole graph. To learn high-quality graph embedding that can be applied to downstream tasks of recommender systems, GNNs need to be extended to take the special properties of recommender system tasks. HyperSoRec (Wang et al., 2021b) proposes that the user-user graph can be approximated as a structure of an $n$-order tree graph. Chami et al. (Chami et al., 2019) show that tree graph embedding in hyperbolic space can achieve better performance than its counterpart in Euclidean space, which may suffer from severe distortion. HyperSoRec develops a hyperbolic mapping layer to map graph embedding in Euclidean space to hyperbolic space. M2GRL (Wang et al., 2020b) proposes to combine homogeneous graphs of multiple views to enhance sparse features and enrich node information. It constructs catalog view, instance view, and shop view, then adopts random walk and SGNS (Levy and Goldberg, 2014) to learn an independent embedding of each view, and finally, maps them to a shared embedding space and aligns them separately. DG-ENN (Guo et al., 2021b) also uses more than one view of homogeneous graphs to learn embeddings jointly. It first learns embeddings from the user-user and item-item homogeneous graphs. Then it leverages the user-item graph as a regularization to ensure the learned embeddings are smooth.
@@ 4.3. Embedding of Bipartite Graphs
The bipartite graph is a unique structure with two node sets, where edges connect nodes from different sets. This is commonly used in recommendation systems (e.g., user-item interaction graphs), represented as $\mathcal{G}=\left(\mathcal{V}_{A}\cup\mathcal{V}_{B},\mathcal{E}\right)$, with $\mathcal{V}_{A}$ and $\mathcal{V}_{B}$ as node sets, and $\mathcal{E}$ denoting the edges. Graph Convolutional Matrix Completion (GC-MC) (Berg et al., 2017) predicts ratings using a graph encoder to aggregate information from neighboring nodes with uniform weight. It aggregates user embeddings from one-hot embeddings based on interacted items, and item embeddings from one-hot vectors of interacting users. A bilinear decoder then reconstructs user-item ratings. However, GC-MC has drawbacks: 1) It employs one-hot encoding, hindering new node processing and becoming inefficient as the graph size grows. 2) GC-MC overlooks node features. STAR-GCN (Zhang et al., 2019) resolves these by end-to-end learning of low-dimensional node embeddings. It simulates embedding new nodes with a masked vector, training the model to reconstruct embeddings. NGCF (Wang et al., 2019b) addresses GC-MC’s second drawback by incorporating node features in its embeddings. LightGCN (He et al., 2020) argues that, as the inputs of the user and item stem from ID embeddings lacking semantic information, there is no need for nonlinear transformations. Therefore, LightGCN introduces a straightforward and efficient aggregation method:
@@
Where $\mathcal{N}_{u}$ and $\mathcal{N}_{i}$ denote the number of neighbor nodes of user $u$ and item $i$. The $\mathbf{h}_{u}^{(l)}$ and $\mathbf{h}_{i}^{(l)}$ denote the node embeddings of user $u$ and item $i$ in layer $l$. Although LightGCN greatly simplifies GCN, it still requires a long training time because multilayer message passing dominates the training of the model. The relatively slow training limits the application of LightGCN in real recommendation scenarios. UltraGCN (Mao et al., 2021) further removes the multi-layer messaging process by directly optimizing the cosine similarity of nodes and neighbors to capture the higher-order collaborative signals between users and items, and it uses a negative sampling strategy to avoid oversmoothing. Recognizing the inherent sparsity within user-item interaction graphs, Collaborative Similarity Embedding (CSE) (Chen et al., 2019) suggests enhancing embeddings by incorporating higher-order similarities among nodes of the same type. CSE introduces two modules, DSEmbed and NSEmbed, to capture user-item similarity and inter-item/inter-user similarity within a user-item bipartite graph. DSEmbed employs an optimization scheme based on rating data to model the proximity between users and items. This is achieved by maximizing the log-likelihood function of positive and negative samples. Conversely, NSEmbed employs random walks and CBOW (Mikolov et al., 2013a) to predict central nodes, enabling the modeling of higher-order neighbor relationships among users or items. Another approach, GE (Xie et al., 2016), proposes joint training to address the sparsity observed in user-POI bipartite graphs. This is achieved by incorporating multiple graph perspectives, where supplementary bipartite graphs (e.g., POI-POI, POI-Region) augment the model’s efficacy by providing enriched information.